# leftmargin=10pt, rightmargin=10pt, vskip=0pt

### v1

@@ 1 Introduction
Figure 1: Overview of BEACON: a: Categorization of the 13 benchmark tasks into classification and regression at both nucleotide and sequence levels. b: Diverse database distinguished by data size and source type. c: Visual representations of tasks across Structure, Function, and Engineering. d: List of baseline models, including naive supervised deep models and advanced RNA language models. e: Metrics for evaluating model performance in classification and regression tasks, tailored to RNA analysis specifics.
@@ 3 Benchmark Tasks
BEACON comprises 13 tasks designed to evaluate RNA models comprehensively, covering structural analysis, functional studies, and engineering applications . The following sections provide detailed information, including data statistics, evaluation metrics, and data sources as shown in Tab 1.
@@ 3.2 Function Prediction
SSP 10,814/1,300/1,305 F1 Multi-label Cls Nucleotide 499/133.8 bpRNA/NAR [18]
@@
VDP 2,155/245/629 MCRMSE Multi-label Reg Nucleotide 130/118.5 OpenVaccine/NMI [71]
@@ 4 Models
We consider three types of baseline models in our benchmark, including naive supervised models, pre-trained language models, and the proposed BEACON-B. We give the details in the following part and summarize them in Tab. 2.
@@ 5.2 Task Pipeline
5.3 Benchmark results on language model
In Tab 3, we report the benchmark results for popular and opensource methods, including literature SOTAs, naive supervised models and existing RNA language models.
ResNet and LSTM are strong naive supervised models. ResNet, which has only been supervised trained downstream, can outperform most if not all language models on some tasks. And LSTM outperforms the other Naive supervised Models on 9 out of 13 tasks, and on many tasks the performance is better by a large margin.
Pre-trained RNA language models have good potential for RNA understanding. It outperforms the previous task-specific SOTA on 8 out of 13 tasks, demonstrating that the additional unsupervised pre-training brings a lot of gain. However, there is still a long way to go on individual tasks, such as contact map prediction and distance map prediction in the structural task, vaccine degradation rate prediction in the engineering task, and CRISPR on- and off-target prediction. Of course, the previous SOTA method used additional features such as secondary structure, but it shows that there is still a lot of room for improvement in the RNA language model.
@@ 5.4 Component Analysis of RNA language models
In Tab 4 , we study the language model component effect from tokenizer and positional encoding.
The single nucleotide tokenizer is a powerful tool for RNA language models. As shown in Tab 11, it achieves the best performance on 11 out of 13 tasks, significantly outperforming other tokenizers. BPE and non-overlapping tokenizers are generally ineffective at the nucleotide level, as they lose precision from overlapping. Similarly, the 6mer approach adds local information before the individual tokens, potentially introducing redundancy. We argue that the single nucleotide tokenizer can learn global information, including surrounding context, through self-attention mechanisms. Thus, using the single nucleotide tokenizer is sufficient, and future work should focus on designing interpretable tokenizers based on single nucleotide units [69, 42].
@@ 5.5 BEACON-B: an Efficient Baseline for RNA Language Models
Based on the above analysis of the different components of the RNA language model, combined with the Tab 4, we use the single nucleotide tokenizer, ALiBi as the positional encoding, and pre-train on human filtered ncRNA sequences from RNACentral [68]. We propose the low-resource and cost-effective BEACON-B (pre-trained on 1026 length seqs) and BEACON-B512 (pre-trained on 512 length seqs and FlashAttn [20]) as a baseline to provide an extremely fast and easy to use open-source pre-training model for subsequent researchers.
Although with very small GPU days (days * GPUs) as the cost of pre-training, BEACON-B can even outperform SOTA pre-trained RNA LMs on some tasks such as contact map prediction and distance map prediction. Compared with other models that also report pre-training resources, BEACON-B and BEACON-B512 can match or even surpass existing RNA language models in one-to-one comparisons on almost half of the tasks listed in Tab 3 and Tab 5, despite being pre-trained with significantly fewer resources. This demonstrates that the insights we obtain from the important components in analysing the RNA language model are vital, and that biological motifs and configurations on limited RNA data can be fully explored by utilising such a combination of components in a good way.
@@ 6 Conclusions
Appendix A Appendix / supplemental material
Optionally include supplemental material (complete proofs, additional experiments and plots) in the appendix. All such materials SHOULD be included in the main submission.
@@ A.1.1 Most of the tasks
Table 6: Configuration settings for most of the tasks training
@@ A.1.2 Vaccine degradation prediction task
Vaccine degradation prediction is trained using the settings shown in the Tab 7.
Table 7: Configuration settings for vaccine degradation prediction training
@@ A.1.3 Nucleotide-nucleotide level tasks
Nucleotide-nucleotide level tasks are trained using the settings shown in the Tab 8. In addition, the representation also follows Fig 2
Table 8: Configuration settings for nucleotide-nucleotide level tasks training
@@
A.2 Methods in Benchmark

### v2

@@ 1 Introduction
Figure 1: Overview of BEACON. a: Categorization of the 13 benchmark tasks into classification and regression at both nucleotide and sequence levels. b: Diverse database distinguished by data size and source type. c: Visual representations of tasks across Structure, Function, and Engineering. d: List of baseline models, including naive supervised deep models and advanced RNA language models. e: Metrics for evaluating model performance in classification and regression tasks, tailored to RNA analysis specifics.
@@ 3 Benchmark Tasks
BEACON comprises 13 tasks designed to evaluate RNA models comprehensively, covering structural analysis, functional studies, and engineering applications . The following sections provide detailed information, including data statistics, evaluation metrics, and data sources as shown in Table 1.
@@ 3.2 Function Prediction
SSP 10,814/1,300/1,305 F1 Multi-label Cls Nucleotide 499/133.8 bpRNA-1m/NC [95]
@@
VDP 2,155/245/629 MCRMSE Multi-label Reg Nucleotide 107/118.5 OpenVaccine/NMI [111]
@@ 4 Models
We consider three types of baseline models in our benchmark, including naive supervised models, pre-trained language models, and the proposed BEACON-B. We give the details in the following part and summarize them in Table 2.
@@ 5.2 Task Pipeline
5.3 Benchmark Results
In Table 3, we report the benchmark results for popular and opensource methods, including literature SOTAs, naive supervised models and existing RNA language models.
ResNet and LSTM are strong naive supervised models. ResNet, which has only been trained on downstream tasks, can outperform most if not all language models on some tasks. LSTM outperforms the other Naive supervised Models on 9 out of 13 tasks, and the performance is better by a large margin on many tasks.
Pre-trained RNA language models have good potential for RNA understanding. It outperforms the previous task-specific SOTA on 8 out of 13 tasks, demonstrating that the additional unsupervised pre-training brings a lot of gains. However, there is still a long way to go on individual tasks, such as contact map prediction and distance map prediction in the structural task, vaccine degradation rate prediction in the engineering task, and CRISPR on- and off-target prediction. Of course, the previous SOTA method used additional features such as secondary structure, but it shows that there is still a lot of room for improvement in the RNA language model.
@@ 5.4 Component Analysis of RNA Language Models
In Table 4 , we study the language model component effect from tokenizer and positional encoding.
The single nucleotide tokenizer is a powerful tool for RNA language models. As shown in Table 17, it achieves the best performance on 11 out of 13 tasks, significantly outperforming other tokenizers. BPE and non-overlapping tokenizers are generally ineffective at the nucleotide level, as they lose precision from overlapping. Similarly, the 6mer approach adds local information before the individual tokens, potentially introducing redundancy. We argue that the single nucleotide tokenizer can learn global information, including surrounding context, through self-attention mechanisms. Thus, using the single nucleotide tokenizer is sufficient, and future work should focus on designing interpretable tokenizers based on single nucleotide units [107, 120, 66].
@@ 5.5 BEACON-B: an Efficient Baseline for RNA Language Models
Based on the above analysis of the different components of the RNA language model, combined with the Table 4, we use the single nucleotide tokenizer, ALiBi as the positional encoding, and pre-train on filtered human ncRNA sequences from RNACentral [104]. We propose the low-resource and cost-effective BEACON-B (pre-trained on 1026 length seqs) and BEACON-B512 (pre-trained on 512 length seqs and FlashAttn [26]) as a baseline to provide an extremely fast and easy to use open-source pre-training model for subsequent researchers.
Although with very small GPU days (days * GPUs) as the cost of pre-training, BEACON-B can even outperform SOTA pre-trained RNA LMs on some tasks such as contact map prediction and distance map prediction. Compared with other models that also report pre-training resources, BEACON-B and BEACON-B512 can match or even surpass existing RNA language models in one-to-one comparisons on almost half of the tasks listed in Table 3 and Table 5, despite being pre-trained with significantly fewer resources. This demonstrates that the insights we obtain from the important components in analysing the RNA language model are vital and that biological motifs and configurations on limited RNA data can be fully explored by utilising such a combination of components in a good way.
@@ 6 Conclusions
Acknowledgments
This work is funded in part by Shanghai Artificial Intelligence Laboratory and supported in part by HKU Startup Fund, HKU Seed Fund for Basic Research, HKU Seed Fund for Translational and Applied Research, HKU IDS research Seed Fund, and HKU Fintech Academy R&D Funding. It is also supported by the Beijing Super Cloud Computing Center serve platform.
Appendix A Appendix
@@ A.1.1 Most of the Tasks
Table 6: Configuration settings for most of the tasks training.
@@ A.1.2 Vaccine Degradation Prediction Task
Vaccine degradation prediction is trained using the settings shown in the Table 7.
Table 7: Configuration settings for vaccine degradation prediction training.
@@ A.1.3 Nucleotide-Nucleotide Level Tasks
Nucleotide-nucleotide level tasks are trained using the settings shown in the Table 8. In addition, the representation also follows Fig 2.
Table 8: Configuration settings for nucleotide-nucleotide level tasks training.
@@
A.2 Detailed Data Preprocessing and More Baselines for Each Task
A.2.1 Secondary Structure Prediction
We followed the preprocessing from the bpRNA-1m dataset [95]. To mitigate sequence redundancy and enhance the diversity of our dataset, we applied an 80% sequence-identity cut-off and restricted the maximum sequence length to below 500 nucleotides, similar to the process described in both referenced articles. This step was crucial in reducing the risk of overfitting and ensuring that our models are trained on genetically distinct samples.
The dataset was strategically split into three parts: a training set (TR0), a validation set (VL0), and a test set (TS0). This separation was carried out randomly to avoid any biases that might affect the evaluation of the model’s performance.
More baselines on SSP are shown in Table 9.
Table 9: Comparison on SSP.
Method Metric (F1%)
Literature SOTA
Contextfold [123] 54.6
CONTRAfold [29] 56.7
E2Efold [17] 13.0
RNAstructure [70] 53.3
RNAsoft [5] 53.5
RNAfold [67] 53.6
Mfold [127] 53.8
LinearFold [50] 55.0
MXfold2 [91] 55.8
Externafold [110] 56.3
SPOT-RNA [95] 61.9
UFold 65.4
Naive Supervised Model
CNN 49.95$\pm$0.82
ResNet 57.26$\pm$3.14
LSTM 58.61$\pm$0.21
Pretrained RNA Language Model
RNA-FM 68.50$\pm$0.54
RNABERT 57.27$\pm$0.30
RNA-MSM 57.98$\pm$0.47
SpliceBERT-H510 64.93$\pm$0.84
SpliceBERT-MS510 43.24$\pm$28.64
SpliceBERT-MS1024 68.26$\pm$0.20
UTR-LM-MRL 59.71$\pm$0.30
UTR-LM-TE&EL 59.57$\pm$0.20
UTRBERT-3mer 60.37$\pm$0.47
UTRBERT-4mer 59.41$\pm$0.45
UTRBERT-5mer 47.92$\pm$8.75
UTRBERT-6mer 38.56$\pm$28.76
Our BEACON-B
BEACON-B 64.18$\pm$0.44
BEACON-B512 58.75$\pm$3.72
A.2.2 Contact Map Prediction
We choose the benchmark datasets used by [103, 12], which are constructed based on a set of nonredundant RNA 3D structures, containing 1786 entries with resolution < 4A initially. Sequences with length < 32nt or > 1000nt or with redundancy over 80% as well as with too few positive points (< 5) are removed. Finally, 221 sequences left are used for training and 80 sequences for testing. Using their pdb files, We then compute the ground truth pairwise tertiary distance, which is defined as the minimal atomic distance of the two bases. Then the pairwise contact is defined as the distance between two bases is less than 8A.
More baselines on CMP are shown in Table 10.
Table 10: Comparison on CMP.
Method Metric (P@L %)
Literature SOTA
PLMC [112] 28
RNAcontact (Seq) [103] 33
RNAcontact (Cov) [103] 59
RNAcontact (SS) [103] 58
RNAcontact 66
Naive Supervised Model
CNN 43.89$\pm$5.53
ResNet 59.59$\pm$0.68
LSTM 40.41$\pm$1.67
Pretrained RNA Language Model
RNA-FM 47.56$\pm$6.73
RNABERT 45.21$\pm$10.87
RNA-MSM 57.26$\pm$15.38
SpliceBERT-H510 45.80$\pm$6.03
SpliceBERT-MS510 52.64$\pm$7.56
SpliceBERT-MS1024 47.32$\pm$3.16
UTR-LM-MRL 45.51$\pm$23.51
UTR-LM-TE&EL 60.32$\pm$7.27
UTRBERT-3mer 51.03$\pm$21.48
UTRBERT-4mer 44.91$\pm$27.56
UTRBERT-5mer 44.71$\pm$7.64
UTRBERT-6mer 51.56$\pm$20.30
Our BEACON-B
BEACON-B 60.81$\pm$1.70
BEACON-B512 61.20$\pm$2.11
A.2.3 Distance Map Prediction
The dataset used for distance map prediction is the same as the contact map prediction. We first compute the pairwise tertiary distance as above. Then we limit the distance from 0 to 20A and regard the value over 20 as 20. Finally, we use 20 to normalize the distance values and obtain a normalized distance map with elements falling into [0, 1].
A.2.4 Structural Score Imputation
We followed the preprocessing procedure as outlined in [40]. The icSHAPE HEK293 cell line raw sequencing data ($\sim$ 300 million clean reads) was downloaded and processed using icSHAPE-pipe to obtain structure profiles for 4,091 transcripts. The full-length transcript structure profiles were then binned into non-overlapping 100-nt fragments. 21,859 fragments were retained with valid structural scores for every nucleotide.
These valid fragments were randomly split into training and test sets in a 7:3 ratio, resulting in 15,085 fragments for training and 6,774 for testing. 569 fragments, with high sequence similarity to any fragment in the training set based on BLAST results, were removed from the test set.
For the training set, 30% of the nucleotides were randomly masked as null, with the ground-truth structural scores of these positions used as training labels. For the test set, the original icSHAPE dataset was downsampled to 100 million clean reads, with the structure profiles to simulate lower sequencing depth recalculated. This process generated missing values, and only fragments with at least one valid structural score were retained, resulting in 3,095 fragments in the final test dataset.
A.2.5 Splice Site Prediction
We followed the preprocessing protocol from [53]. The GENCODE V24lift37 gene annotation table was downloaded from the UCSC table browser, and 20,287 protein-coding gene annotations were extracted. For genes with multiple isoforms, the principal transcript was selected. Genes lacking splice junctions were removed. The remaining genes were split into training and test sets: chromosomes 2, 4, 6, 8, 10-22, X, and Y were used for training, with 1/9 reserved for early stopping. For testing, genes from chromosomes 1, 3, 5, 7, and 9 without paralogs were selected.
For each gene, mRNA transcript sequences from the canonical transcription start to end sites were extracted using the hg19/GRCh37 assembly. Sequences were zero-padded to a multiple of 5,000 nucleotides and split into blocks of length 5,000. Then, the center 100 nucleotides of each block were extracted as the input sequence, and the corresponding splice site label was extracted as the output. (non-splice sites [1, 0, 0], splice acceptors [0, 1, 0], and splice donors [0, 0, 1])
A.2.6 APA Isoform Prediction
The preprocessing for IPA Isoform began with filtering the raw sequencing reads from all MPRAs [94] to ensure high-quality, full-length RNA reads. These reads were clustered based on the randomized region upstream of the proximal polyadenylation site (pPAS), creating a dictionary of sequence variants for each library. To expand the dictionary, the plasmid library was sequenced to include members that did not express a distal isoform, and RNA reads were mapped to their respective dictionary entries by matching the upstream region with the shortest Hamming distance.
For each mapped read, the polyadenylation cleavage site was identified by scanning for the Poly-A tail. The cleavage positions were stored as vectors associated with each sequence variant, including a special position for reads mapping to non-random distal sites. The resulting dataset consisted of a dictionary of unique sequence variants with associated cleavage-position count vectors, which then underwent a final filtering step. This step selected high-confidence variants by removing sequences supported by fewer than 10-20 unique UMI RNA reads or sequences with over 75% A-nucleotides in a 12-20 bp region to avoid internal priming artifacts.
Data from 12 random 3’ UTR libraries were processed, with 9 used for training and 3 held out (the 3 held-out libraries were not used in the current analysis). To ensure balanced representation in the test set, the sequences from each library were individually shuffled based on the read count, then combined using a round-robin order, placing one sequence from each library after another in descending order of read count. This process ensured that the test set contained an equal number of high-read count sequences from each library. The remaining sequences were placed at the beginning of the combined library, and the training set was further shuffled. This approach ensured a balanced, high-quality dataset for training, validation, and testing. To maintain balanced and high-quality data for benchmarking, the top 10% of high-read count sequences were selected, and the most highly expressed sequences were further chosen for testing.
A.2.7 Non-coding RNA Function Classification
We followed nRC [34] and sourced non-coding RNA sequences from the Rfam database, recognized for its comprehensive collection of manually curated RNA sequences. We specifically selected 13 diverse ncRNA classes including miRNA, 5S rRNA, 5.8S rRNA, ribozymes, CD-box, HACA-box, scaRNA, tRNA, Intron gpI, Intron gpII, IRES, leader and riboswitch. Utilizing the CD-HIT tool [35], we reduced sequence redundancy to 20%, ensuring a representative yet manageable dataset size. This process allowed us to include a wide array of RNA types while controlling data complexity.
For each selected class, we systematically assembled a training-validation dataset comprising 500 sequences per class, except for the IRES class where only 320 sequences were available, leading to a total of 6320 ncRNA sequences. We subsequently divided it into two groups: a validation set comprising 650 sequences, with 50 from each class, and a training set consisting of the remaining 5670 sequences. The balancing was further refined in our test dataset, which contains an equal number of sequences from each class, totaling 2600 sequences, ensuring that our model evaluation would not be biased by uneven class representation.
More baselines on ncRNA are shown in Table 11.
Table 11: Comparison on ncRNA.
Method Metric (ACC %)
Literature SOTA
EDeN [71] 67
nRC [34] 81.81
RNAGCN 85.73
Naive Supervised Model
CNN 88.62$\pm$0.71
ResNet 88.33$\pm$1.22
LSTM 88.78$\pm$0.10
Pretrained RNA Language Model
RNA-FM 96.81$\pm$0.061
RNABERT 68.95$\pm$7.285
RNA-MSM 84.85$\pm$0.266
SpliceBERT-H510 95.92$\pm$0.666
SpliceBERT-MS510 95.87$\pm$0.364
SpliceBERT-MS1024 96.05$\pm$0.777
UTR-LM-MRL 89.97$\pm$0.617
UTR-LM-TE&EL 81.33$\pm$8.551
UTRBERT-3mer 92.88$\pm$0.379
UTRBERT-4mer 94.32$\pm$0.946
UTRBERT-5mer 93.04$\pm$0.367
UTRBERT-6mer 93.12$\pm$0.168
Our BEACON-B
BEACON-B 94.63$\pm$0.16
BEACON-B512 94.99$\pm$0.21
A.2.8 Modification Prediction
We followed MultiRM [97] and compiled a comprehensive dataset comprising 20 epi-transcriptome profiles derived from 15 different base-resolution technologies, addressing 12 distinct RNA modifications such as m6A, m1A, m5C, among others. A key focus was given to constructing a reliable set of negative control data, which was crucial for our predictors’ accuracy. To achieve this, negative sites (non-modified nucleotides) were carefully selected from the unmodified bases within the same transcripts that contained the positive modification sites, ensuring an authentic comparison baseline.
Moreover, the dataset was divided into three distinct subsets: training, validation, and testing. The training set was purposefully left unbalanced to mirror the natural prevalence differences among the RNA modification types, which introduces realistic challenges in model training akin to real-world scenarios. Conversely, the validation and test sets were meticulously balanced—each set containing 150 and 50 samples, respectively, across all modification types—to ensure fairness in model evaluation and performance metrics.
A.2.9 Mean Ribosome Loading
For 5’UTR processing, we primarily employed a large-scale synthetic Human 5’UTR library [88], which comprised 83,919 sequences of varying lengths, as articulated in the first referenced article. The validation set was meticulously crafted by evenly sampling 7600 sequences across these different lengths to ensure fair representation and robust generalizability testing. The remaining sequences were utilized for training purposes. To further enhance our model’s performance assessment, we incorporated an additional dataset of 7600 real human 5’UTRs, mirroring the length distribution provided by the synthetic library. This dual-dataset strategy not only allowed for a thorough evaluation of the model’s predictive accuracy but also its ability to generalize across synthetic and real-world data.
Additionally, as described in the second referenced article, the pivotal role of the 5’UTR sequence in translation efficiency was explored through rigorous experiments utilizing data from Massively Parallel Reporter Assays (MPRAs), which included a substantial library of 280,000 gene sequences.
A.2.10 Vaccine Degradation Prediction
We followed OpenVaccine [111] to collect 2400 sequences specifically for training purposes, 629 sequences were made available for public testing during the competition, and the remainder were reserved for private scoring. Post-competition, all these sequences are now available with comprehensive labels detailing various degradation rates under multiple experimental conditions, such as high pH and high temperature, both with and without magnesium.
To ensure the integrity of the public leaderboard and prevent biases, we implemented rigorous filtering criteria on the 629 sequences used for public testing. These criteria included setting a minimum value threshold across all experimental conditions and ensuring a mean signal-to-noise ratio above a specified level.
We also provide the test data of the original private leaderboard although we do not use it in the paper. For the private leaderboard data, which involves the most critical evaluation, we included measurements from an additional 3005 RNA sequences, which are slightly longer. These were processed to include data for the first 91 bases, carefully excluding the final bases to align with experimental limitations.
A.2.11 Programmable RNA Switches
We followed the data generation pipeline from [7]. A toehold-switch library based on 244,000 putative trigger sequences was designed and synthesized, which covered the complete genomes of 23 pathogenic viruses, the entire coding regions of 906 human transcription factors, and approximately 10,000 random sequences. The synthesized oligo pool was used to generate two construct libraries for ON and OFF states, which were transformed into BL21 E. coli. The OFF library contained toehold-switch constructs without triggers, while the ON library contained the same toeholds with complementary triggers fused to their corresponding switches.
These libraries were sorted using fluorescence-activated cell sorting (FACS) into four bins, and the variants in each bin were quantified using next-generation sequencing (NGS) to determine their fluorescence distributions. After quality control, the toehold-switch library included 109,067 ON-state measurements, 163,967 OFF-state measurements, and 91,534 ON/OFF paired ratios, where both states were characterized for a given switch. The ON and OFF data were normalized from 0 to 1, resulting in ON/OFF ratios normalized from -1 to 1. Following  [7], a quality control process was applied to remove artifacts and ensure the reliability of the data. There are 5 levels of quality control (QC1, QC2, QC3, QC4, and QC5), with QC1 being the lowest quality and QC5 being the highest. All datasets at QC levels above QC2 are used for training with Q5 left for testing.
Table 12: Comparison on CRI-On.
Method Metric (SC %)
Literature SOTA
CHOPCHOP [60] 0.9
DeepCRISPR [19] 26.2
WU-CRISPR [114] 26.8
sgRNA Scorer [11] 30.5
CRISPR MultiTargeter [77] 32.5
E-CRISP [45] 33.6
sgRNA Designer [30] 41.8
SSC 44.1
Naive Supervised Model
CNN 29.69$\pm$2.52
ResNet 28.55$\pm$2.42
LSTM 26.83$\pm$1.32
Pretrained RNA Language Model
RNA-FM 31.62$\pm$1.16
RNABERT 29.77$\pm$3.98
RNA-MSM 34.92$\pm$1.99
SpliceBERT-H510 26.61$\pm$1.30
SpliceBERT-MS510 27.13$\pm$0.27
SpliceBERT-MS1024 27.59$\pm$4.61
UTR-LM-MRL 28.49$\pm$1.37
UTR-LM-TE&EL 32.49$\pm$4.14
UTRBERT-3mer 29.92$\pm$1.95
UTRBERT-4mer 23.20$\pm$1.10
UTRBERT-5mer 25.74$\pm$0.00
UTRBERT-6mer 28.60$\pm$1.55
Our BEACON-B
BEACON-B 26.01$\pm$1.81
BEACON-B512 28.17$\pm$1.81
A.2.12 CRISPR On-Target Prediction
For the processing of the on-target sgRNA dataset, we sourced experimentally validated sgRNAs targeting approximately 1,071 genes across four distinct cell lines [19], namely hct116, hek293t, hela, and hl60. To ensure the integrity of our dataset, we selected hl60 cell line data and removed redundant entries and restricted our dataset to sgRNAs with direct experimental validation of knockout efficacy, quantified as the log-fold change.
Furthermore, for a balanced and normalized dataset conducive to regression analysis, we employed a collaborative filtering-based normalization approach  [19], akin to methodologies used in user-item recommendation systems. We constructed an efficacy matrix where rows represented experiments and columns represented sgRNAs. The knockout efficacy was normalized by calculating the mean values across the rows, columns, and the entire matrix, and then adjusting the sgRNA efficacy scores by subtracting these mean values and dividing by the number of mean types considered. This normalization process ensures that our dataset remains unbiased and reflective of true knockout efficacies across various experimental setups, thereby enhancing the generalizability and accuracy of our predictive models.
More baselines on CRI-On are shown in Table 12.
Table 13: Comparison on CRI-Off.
Method Metric (SC %)
Literature SOTA
CFD [30] 10.3
MIT [47] 8.0
CCTop [98] 5.2
CROP-IT [96] 4.2
DeepCRISPR 12.6
Naive Supervised Model
CNN 11.40$\pm$0.10
ResNet 11.50$\pm$0.22
LSTM 8.60$\pm$0.13
Pretrained RNA Language Model
RNA-FM 2.49$\pm$1.56
RNABERT 4.27$\pm$1.05
RNA-MSM 3.85$\pm$0.99
SpliceBERT-H510 4.00$\pm$1.13
SpliceBERT-MS510 3.49$\pm$2.12
SpliceBERT-MS1024 5.00$\pm$0.71
UTR-LM-MRL 4.28$\pm$0.15
UTR-LM-TE&EL 2.91$\pm$1.18
UTRBERT-3mer 4.48$\pm$1.12
UTRBERT-4mer 3.11$\pm$1.10
UTRBERT-5mer 3.93$\pm$0.24
UTRBERT-6mer 4.90$\pm$0.57
Our BEACON-B
BEACON-B 4.42$\pm$0.33
BEACON-B512 3.82$\pm$1.04
A.2.13 CRISPR Off-Target Prediction
We followed the dataset from DeepCRISPR [19] including off-target profiles from two distinct cell types: 293-related cell lines and K562 cells, encompassing 30 sgRNAs in total. Using the bowtie2 tool  [19], we used K562 cell data and identified approximately 160,000 potential off-target loci across the genome, allowing for up to six nucleotide mismatches per sgRNA. This resulted in a highly unbalanced dataset, with varying numbers of loci associated with each level of mismatch, from one to six.
To address the imbalance and refine our dataset for the regression models, we labeled the off-target sites and normalized them according to the indel frequency detected by various genome-wide off-target detection techniques. This normalization process helped to mitigate the skewness introduced by the uneven distribution of loci, ensuring that our model could generalize effectively across different genomic backgrounds and detection assays.
More baselines on CRI-Off are shown in Table 13.
A.3 Methods in Benchmark